Circular No. 7316

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ga IN NGC 2442 S. Woodings, R. Martin, and A. Williams, Perth Observatory, report

their discovery of an apparent supernova (R ∼ 18) on Nov. 19.76 UT in the course of the Perth Astronomy Research Group’s automated supernova search with the Perth-Lowell 0.61-m reflector. The last previous image of the field was taken on 1999 Apr. 30, with no object visible at the position of SN 1999ga to mag ∼ 19. J. Biggs and A. Verveer provide the following position for SN 1999ga from two unfiltered CCD images obtained on Nov. 22.54 with the 0.25-m Mike Candy Telescope at Perth Observatory: α = 7^{h}36^{m}16^{s}.70 ± 0^{s}.03, δ = −69^{o}33′21′′.8 ± 0′′.4 (equinox 2000.0; GSC reference stars), which is about 38′′ west and 91′′ south of the center of NGC 2442; the magnitude of the new star was estimated as 18.5 ± 1.

SUPERNOVA 1999gb IN NGC 2532 A. Friedman and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova in NGC 2532. SN 1999gb was discovered and confirmed on unfiltered images taken on Nov. 22.4 (mag ∼ 16.1) and 23.4 UT (mag ∼ 15.9). SN 1999gb is located at α = 8^{h}10^{m}13^{s}.70, δ = +33^{o}57′29′′.8 (equinox 2000.0), which is 18′′.6 west and 6′′.2 north of the nucleus of NGC 2532. At the request of Li, an image taken by T. Puckett with the Puckett Observatory 0.60-m automated supernova patrol telescope on Nov. 23.4 also showed the new object. A KAIT image of the same field on Nov. 14.4 (limiting mag ∼ 18.0) showed nothing at the position of SN 1999gb.

COMET C/1999 J3 (LINEAR) Visual m_{1} estimates: Oct. 31.09 UT, 8.3 (A. Amorim, Florianopolis,

Brazil, 0.14-m reflector); Nov. 3.80, 8.4 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars); 11.52, 9.6 (M. Mattiazzo, Wallaroo, S. Australia, 0.20-m refl.); 20.40, 9.5 (S. T. Rae, Te Awamutu, New Zealand, 0.25-m refl.).

HD 209458 Corrigendum. On IAUC 7315, line 26, for G. A. Druckier, read G. A.

Drukier,

1999 November 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green